The probabilities that a furnace station will supply melted products to 0,1,2,3,4, 5 or more temperature controlled trucks during a certain 30-minute period are 0.03, 0.18, 0.24, 0.28, 0.10 and 0.17. Find the probability that in this 30-minute period (a) more than two trucks received melted products (b) at most 4 trucks received melted products and (c) 4 or more trucks received melted products. 


1
Expert's answer
2021-10-05T11:24:17-0400

"P(X=0) = 0.03 \\

P(X=1) = 0.18 \\

P(X=2) = 0.24 \\

P(X=3) = 0.28 \\

P(X=4) = 0.10 \\

P(X≥5) = 0.17"

(a)

"P(X>2) = 1 -P(X≤2) \\

= 1 -[P(X=0) + P(X=1) + P(X=2)] \\

= 1 -[0.03+0.18+0.24] \\

= 1 -0.45 \\

= 0.55"

(b)

"P(X≤4) = P(X=0) +P(X=1) + P(X=2) + P(X=3) + P(X=4) \\

= 0.03 + 0.18 + 0.24 +0.28 +0.1 \\

= 0.83"

(c)

"P(X≥4) = P(X=4) +P(X=5) \\

= 0.10 + 0.17 \\

= 0.27"